OpenStudy (anonymous):

The resistance of a conductor is reduced to half its initial value. In doing so the heating effects in the conductor will become : a) Half b) One-fourth c) Four times d) Double

OpenStudy (anonymous):

@Fifciol

OpenStudy (anonymous):

@Festinger

OpenStudy (festinger):

The heating effect of a conductor is given by: \[P=VI\] Assume the conductor is ohmic and thus obeys ohm's law: \[V=IR\] Putting this in the first question you can see that \[P=I^{2}R\] Before making any changes, \[V=I_{1}R_{1}\] and so power is \[P_{1}=\frac{V^{2}}{R_{1}}\] But after we reduce the resistance, current will be another value, following ohm's law: \[V=I_{2}R_{2}\] But \[R_{2}=\frac{1}{2}R_{1}\] and so \[V=\frac{I_{2}R_{1}}{2}\] And thus \[I_{2}=\frac{2V}{R_{1}}\] and power is then: \[P_{2}=I^{2}R_{2}=\frac{4V^{2}R_{1}}{2R_{1}^{2}}=\frac{2V^{2}}{R_{1}}\] Comparing with the first result the heating effect doubles.

OpenStudy (imtiaz7):

the heat effect increases because voltage is directly proportional to current as there it there is some heat but when the resistance of a conductor is reduced to half its initial value then the heat become d) doubled for example: the high voltage fire the materials.

OpenStudy (anonymous):

power dissipiated as heat=v^2/r.. if r is halved.. heat produced doubles

OpenStudy (ybarrap):

\[P _{old} = I _{old} ^{2}R _{old}\] \[I _{new}=\frac{ V }{ R _{old}/2 }=\frac{ 2V }{ R _{old} }=2I _{old}\] \[P _{new}=I _{new}^{2}R _{new}=2^{^2} I _{old }^{2}R _{old}/2=\frac{ 4 }{ 2 } I _{old }^{2}R _{old}=2P _{old}\] It doubles.

OpenStudy (anonymous):

thank you so much guys....for helping me @Festinger @Imtiaz7 @mr.singh @ybarrap